OpenStudy (anonymous):

2/3 (divided by) 0.25= In fraction form????

OpenStudy (rulnick):

2/3 / 0.25 = 2/3 / (1/4) = 2/3 * 4/1 = ... Get it from here?

OpenStudy (rulnick):

Alternatively, you may view division by 0.25 (or 1/4) as multiplication by 4 (or 4/1). So 2/3 times 4 is ... Get it now?

OpenStudy (anonymous):

ok.....yeah I get it.. But my teacher writes it differently. Aren't you supoed to times it?

OpenStudy (rulnick):

Well, yes. In the first case you have fraction multiplication, 2/3 * 4/1 = (2*4) / (3*1) = 8/3. In the second case it's just 4 * 2/3 = 8/3.